### v1

# PALADIN : Robust Neural Fingerprinting for Text-to-Image Diffusion Models

Abstract
The risk of misusing text-to-image generative models for malicious uses, especially due to the open-source development of such models, has become a serious concern. As a risk mitigation strategy, attributing generative models with neural fingerprinting is emerging as a popular technique. There has been a plethora of recent work that aim for addressing neural fingerprinting. A trade-off between the attribution accuracy and generation quality of such models has been studied extensively. None of the existing methods yet achieved $100\%$ attribution accuracy. However, any model with less than perfect accuracy is practically non-deployable. In this work, we propose an accurate method to incorporate neural fingerprinting for text-to-image diffusion models leveraging the concepts of cyclic error correcting codes from the literature of coding theory.
1 Introduction
The current state-of-the-art text-to-image generative models have shown prodigious capabilities in generating high-quality and diverse images from natural language text prompts. This has enabled the widespread creation and manipulation of photo realistic images, adhering to precise textual prompts. This has led to many image editing tools [2] [4] [11] that are becoming creative resources for artists, designers, and the public. Although this is a great leap forward for generative modeling, it raises concerns about the authenticity, veracity, validity, and provenance of the generated images [5]. The increasing realism of such generated images poses a serious threat to both individual and social interests. A convincing and widely shared misinformation could disrupt democratic discussion, influence elections, or threaten national security. A viable solution to this problem is to create accountability for the generated image by integrating user information in the generated artifacts.
A line of work in this area has emerged that analyze design principles of fingerprinting strategies in text-to-image diffusion models and the trade-off between the attribution accuracy and generated image quality. However, none of such methods is able to achieve 100% accuracy. For example, an attribution accuracy of 0.99% means the aforesaid method will still predict $100,000$ user incorrectly for a database containing 1 billion users. Unless a method produces absolute zero incorrect predictions, such near-perfect methods can not be deployed in the real world.
Figure 1: PALADIN: a robust and model-agnostic strategy that elevates a near-perfect latent diffusion based fingerprinting strategy to 100% attribution accuracy utilizing binary cyclic error correction codes.
WOUAF
Generated Image (1)Refer to caption Pixel wise Difference 2x (1) Refer to caption Generated Image (2)Refer to caption Pixel wise Difference 2x (2)Refer to caption
PALADIN
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: A qualitative comparison of PALADIN with WOUAF on COCO dataset. Pixelwise difference being measured w.r.t original image and amplified by 2x. PALADIN shows superior accuracy with no significant image residual. Prompts used 1) ’A city street at night with a tall tower at the end of a road.’ and 2) ’This cake is being doused in liquid chocolate’.
Named after legendary nights and defender of a noble cause, we introduce the PALADIN (Perfect user Attribution for LAtent DIffusioN) framework. PALADIN leverages cyclic error correcting codes to address the robust neural fingerprinting for latent diffusion models. It is capable of elevating a near-perfect fingerprinting in latent diffusion models to be a perfect one. PALADIN is capable of handling image post-processing and detect responsible user for generated artifacts with $100\%$ confidence and also appropriately flagging non-identifiable fingerprints.
Our implementation is based atop a state-of-the-art method, WOUAF [7]. However, the core principle can be applied to other models as well. To summarize, below is the list of our contributions.
  • First, we propose to improve both the image generation quality and attribution accuracy on top of our strongest baseline, WOUAF, with a combination improvised module architecture and loss functions.
  • We propose a framework, PALADIN, for accurate fingerprinting to elevate a near-perfect model to a perfect one utilizing cyclic error correction codes.
  • Finally, we introduce FER (Fingerprint Error Rate), an informative and accurate metric to measure the effectiveness of such models. We show the efficacy of the model by extensive experiments on standard benchmarks.
To the best of our knowledge, this is the first method that achieves 100% attribution accuracy and the lowest FER for the text-to-image diffusion model.
2 Related Work
Yu et al. [13] proposed a weight modulation technique that modulates the generator’s weights to embed user information in the generated artifacts. By incorporating fingerprints in the generator parameter rather than generator input facilitates in training a generic model that can be instantiated with unique fingerprint catering to a large user base. Despite these advancements, the above method caters to GAN based models and their suitability and effectiveness for diffusion based models remained unexplored. Zhang et al. [15] proposed a fingerprinting technique for latent diffusion models that embed fingerprinting in the Fourier space of the latent vector. However, this method is restricted to a single watermark per training cycle. Fernandez et al. [3] proposed a fingerprint embedding technique which embeds user information into generated artifacts by finetuning the latent decoder. A pretrained fingerprint encoder-decoder network is used to finetune the decoder to embed user information in the generated artifacts. However, this approach requires one train cycle per user, necessitating high compute and time. Kim et al. [7] proposed a weight modulation technique in which weights of the Stable Diffusion (SD) are modulated to embed user information in the generated artifacts. This facilitates the generation and distribution of fingerprint decoders, capable of handling multiple embeddings, with a single training cycle. Although this method achieves satisfactory results against attacks such as image augmentation at the cost of robustness and probabilistic attribution accuracy rendering it un-deployable in real-world scenarios for tracking the malicious user responsible for generating artifacts.
One of the recent state-of-the-art method ,WOUAF [7], introduced the concept of embedding the watermarks in all the images they generate by fine-tuning of the decoder of Latent Diffusion Models. WOUAF [7] is based on StyleGAN [6] weight modulation technique to embed user information in the generated artifacts. These watermarks are more robust than previously existing models could generate, invisible to the human eye and could be employed to detect generated images and identify the user that generated it, with high performance. In ours, we make use of recent progress in fingerprinting in diffusion models and enhance their effectiveness with the power of error correction capability. PALADIN can broadly be classified as distributor-oriented fingerprinting, in which the model distributor assigns and embeds a unique traceable fingerprint to each user who downloads the model. In particular, we build atop WOUAF. However, the core principle can be applied to other models as well.
3 Approach
     Methods      Err. Detection      Bit Acc.($\uparrow$)      FER ($\downarrow$)
     WOUAF      ✗      0.9974      0.0700*
     PALADIN $\psi$      ✗      0.9999      0.0020*
     PALADIN $\phi$      ✔      1.0000      0.0004
Table 1: A comparison of fingerprint accuracy and fingerprint error rate (FER). Experiments were conducted on 512 x 512 images on MS-COCO validation dataset, with 32bit fingerprint. Error detection capability is the ability of fingerprinting decoder block to report corrupted fingerprint. Bit accuracy is the proportion of correctly decoded bits to the total bits. Fingerprint error rate is the proportion of corrupted fingerprint to the total number of fingerprints.
*fingerprint corruption are not reported, hence, calculated in test bench with reference fingerprint.
3.1 Preliminaries
Our work is primarily designed for diffusion based models but should be applicable for GAN based models as well. The primary focus is on open source Stable Diffusion (SD) model, that consists of $Encoder$ $\mathcal{E}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}};z=\mathcal{E}(x)$ which maps pixel space to a latent space, U-Net based diffusion model $\epsilon_{\theta}$, cross attention mechanism for textual input and $Decoder$ $\mathcal{D}:\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}^{d_{x}};x=\mathcal{D}(z)$ which maps latent space to pixel space.
3.2 Overview
Our primary objective is to fine-tune SD’s decoder such that generated artifacts contain uniquely identifiable user information without significant change in image quality and perception, this is achieved using weight modulation technique [14, 6]. Our secondary objective is to train a deciphering network capable of extracting ciphered fingerprint from the generated artifacts. The overview of the neural fingerprinting architecture is shown in Fig. 1. A pretrained SD is used as the base, only the $decoder$ $\mathcal{D}$ is fine-tuned keeping the reset of the components untouched.
3.3 Fingerprint encoder
A binary fingerprint $\phi\in\{0,1\}^{d_{\phi}}\sim Ber(0.5)^{d_{\phi}}$, of length $d_{\phi}$, be uniquely identifiable user information that needs to be embedded in the generated artifacts. We utilize a fingerprint encoder $\mathcal{F_{E}}:\mathbb{R}^{d_{\phi}}\rightarrow\mathbb{R}^{d_{\upsilon}};\upsilon=\mathcal{F_{E}}(\phi)$ that is responsible for generating the ciphered fingerprint $\upsilon$. The fingerprint encoder block consists of 1) The BCH encoder $\mathcal{E_{BCH}}:\mathbb{R}^{d_{\phi}}\rightarrow\mathbb{R}^{d_{\psi}};\psi=\mathcal{E_{BCH}}(\phi)$ based on the Bose–Chaudhuri–Hocquenghem code [1], a class of cyclic error-correcting codes capable of detecting and correcting multi-bit errors. This block is responsible for generating unique encoded fingerprint $\psi$ from the user fingerprint $\phi$. 2) The cipher network block $\mathcal{E_{\psi}}:\mathbb{R}^{d_{\psi}}\rightarrow\mathbb{R}^{d_{\upsilon}};\upsilon=\mathcal{E_{\psi}}(\psi)$ that generates the ciphered fingerprint $\upsilon$.
The weight modulation of the decoder $\mathcal{D}$ is achieved using StyleGAN [10] based weight modulation, for the layer $l$, ${W_{l}}^{\prime}=\mathcal{A}_{l}({\upsilon})*W_{l}$, here $\mathcal{A}_{l}:\mathbb{R}^{d_{\upsilon}}\rightarrow\mathbb{R}^{d_{l}}$ is the $l^{th}$ affine transformation realized using the fully connected network responsible for the dimensionality matching. The weight modulated SD decoder $\mathcal{D}:\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}^{d_{W}\times d_{H}\times d_{C}};\tilde{x_{\psi}}=\mathcal{D}(z)$
Methods SSIM($\uparrow$) PSNR($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$)
WOUAF 0.9206 28.2097 0.0731 7.6668
PALADIN 0.9534 30.4530 0.0559 6.2240
Table 2: A study and comparison of image quality for a 32 bit fingerprint being embedded in an image. All the metrics are calculated w.r.t output of unmodified SD.
Method SSIM($\uparrow$) PSNR($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$)
SD 0.8860 28.4201 0.1047 09.4744
WOUAF 0.8746 27.1743 0.1208 13.6315
PALADIN 0.9025 28.4982 0.1108 14.5790
Table 3: A study and comparison of image quality for a 32 bit fingerprint being embedded in an image. All the metrics are calculated w.r.t original dataset image.
3.4 Fingerprint decoder
The embedded fingerprint is extracted from the artifact using the fingerprint decoder $\mathcal{F_{D}}:\mathbb{R}^{d_{W}\times d_{H}\times d_{C}}\rightarrow\mathbb{R}^{d_{\phi}}$. The fingerprint decoder consists of a decipher network $\mathcal{D_{\psi}}:\mathbb{R}^{d_{W}\times d_{H}\times d_{C}}\rightarrow\mathbb{R}^{d_{\psi}};\psi=\mathcal{D_{\psi}}(\tilde{x_{\psi}})$ that is responsible for extracting the embedded fingerprint from the artifacts, and the BCH decoder $\mathcal{D_{BCH}}:\mathbb{R}^{d_{\psi}}\rightarrow\mathbb{R}^{\phi};\phi=\mathcal{D_{BCH}}(\psi)$ based on Bose–Chaudhuri–Hocquenghem code [1] capable of detecting, correcting, and reporting fingerprint corruption.
Method Brightness Contrast Saturation Sharpness H Flip G Noise Crop JPEG Comp.
(1.35) (1.35) (1.35) (1.35) $(\sigma=0.1)$ (20%) (qual. = 60)
WOUAF 0.99872 0.99884 0.99887 0.99281 0.99744 0.99910 0.99754 0.99004
PALADIN 1.00000 1.00000 1.00000 1.00000 1.00000 1.00000 1.00000 1.00000
(a) Table shows the fingerprint accuracy against different image post processing.
Method Brightness Contrast Saturation Sharpness H Flip G Noise Crop JPEG Comp.
(1.35) (1.35) (1.35) (1.35) $(\sigma=0.1)$ (20%) (qual. = 60)
WOUAF* 0.03622 0.03302 0.03342 0.08421 0.07045 0.02301 0.06745 0.22357
PALADIN 0.00080 0.00060 0.00100 0.00800 0.00060 0.00120 0.01120 0.01241
(b) Table shows the fingerprint error rate against different image post processing. *WOUAF does not report fingerprint corruption, all the data calculated are on testbench with reference fingerprint.
Table 4: Accuracy and FER readings against varied post-processing operations.
3.5 Loss function
As part of the training, our primary objective is to maximize the fingerprint decoding capability of the deciphering network. This is achieved through the BCE loss function denoted as $L_{\psi}$ The loss between original fingerprint $\psi$ and extracted fingerprint $\tilde{\psi}$ is
$L_{\psi}=-\sum_{i=1}^{d_{\psi}}\psi_{i}ln(\sigma(\tilde{\psi}_{i}))+(1-\psi_{i})ln(1-\sigma(\tilde{\psi}_{i}))$ (1)
Our secondary objective is to ensure negligible effect of fingerprint on the generated artifacts. To ensure this, we employ three loss components, the first LPIPS perceptual loss [16] $L_{LPIPS}$ responsible for perceptual similarity between original and generated artifacts, second SSIM loss $L_{SSIM}$ responsible for maintaining structural integrity, and third MSE loss $L_{MSE}$. Thus, secondary loss function $L_{x}$ is given by
$\begin{split}L_{x}=\lambda_{1}L_{LPIPS}(x,\hat{x_{\psi}})+\lambda_{2}L_{SSIM}(x,\hat{x_{\psi}})+\\ \lambda_{3}L_{MSE}(x,\hat{x_{\psi}})\end{split}$ (2)
where $\lambda_{1}=\lambda_{2}=\lambda_{3}=1$. The total loss function is given by
$L=L_{\psi}(\psi,\tilde{\psi})+L_{x}(x,\hat{x_{\psi}})$ (3)
Ciphering network $\mathcal{E_{\psi}}$, SD decoder $\mathcal{D}$ and the deciphering network $\mathcal{D_{\psi}}$ are jointly optimized to ensure robust fingerprinting.
4 Experiments
4.1 Experimental Setup
Dataset
We fine-tune SD decoder $\mathcal{D}$ and decipher network $\mathcal{D_{\psi}}$ to generate and decode a 512p resolution image based on MS-COCO [8] dataset, incorporating the Karpathy split.
Experimental Setup
We adopt Stable Diffusion 2.0 for fine-tuning with guidance scale 7.5 and 20 diffusion steps. Our weight modulation of SD Decoder $\mathcal{D}$ is inspired by StyleGAN2-ADA [6], cipher network $\mathcal{E_{\psi}}$ is implemented using two chained fully connected network. The deciphering network $\mathcal{D_{\psi}}$ is implemented using ConvNext [9] with the last classification layer replaced by Layer Norm and FC layers, whereas WOUAF used Resnet50 and FC layer. BCH encoder $\mathcal{E_{BCH}}$ and decoder $\mathcal{D_{BCH}}$ is configured to have a codeword length of 63 bits, message length of up to 39 bits, in this configuration it can detect and correct up to 4 bit errors in $\psi$.
4.2 Evaluation
We evaluated the fine-tuned SD’s decoder with the following metrics: $\mathcal{D}$ using PSNR, SSIM [12] and LPIPS [16]. The accuracy of fingerprint decoding is measured by
$Acc=\frac{1}{d_{\phi}}\sum_{i=0}^{d_{\phi}}\mathcal{1}(\phi_{i}=\hat{\phi_{i}})$ (4)
We introduce and incorporate a fingerprint error rate (FER) metric that evaluates the number of corrupted fingerprints post decoding, formulated as
$\textbf{FER}=\frac{Number\ of\ corrupted\ fingerprint}{Total\ number\ of\ fingerprint}$ (5)
Fingerprint Accuracy
As demonstrated in Table 1, PALADIN outperforms WOUAF (our strongest baseline) significantly and produces robust and superior performance as per all metrics. Even pre-error correction, PALADIN $\psi$’s bit accuracy is significantly higher than WOUAF. A key feature of PALADIN is the ability to detect and report fingerprint corruption.
Image Quality
As shown in Table 2, the quality of generated images with PALADIN is closer to the original SD’s compared to WOUAF as measured by by SSIM, PSNR, LPIPS, and FID. Table 3 shows image quality assessment for PALADIN compared to Stable Diffusion and WOUAF. PALADIN performs better than WOUAF in terms of SSIM and PSNR, while improves upon SD in terms of LPIPS and FID. Figure 2 shows qualitative results which highlight PALADIN’s ability to reliably incorporate fingerprinting without degrading image generation quality.
[Plot showing bit accurary for different brightness scale] [Plot showing FER for different brightness scale]
[Plot showing bit accurary for different contrast scale] [Plot showing FER for different contrast scale]
[Plot showing bit accurary for different sharpness scale] [Plot showing FER for different sharpness scale]
[Plot showing bit accurary for different saturation scale] [Plot showing FER for different saturation scale]
Figure 3: Plots showing accuracy and FER for varying degree of post processing on images.
[Plot showing bit accurary for different jpeg compression scale] [Plot showing FER for different jpeg compression scale]
Figure 4: Plots showing accuracy and FER for varying degree of jpeg compression on images.
4.3 Resilience against image post-processing
One of the most common techniques to corrupt or erase fingerprinting is image post-processing techniques, in that a malicious user attempts to obscure fingerprint by employing one or more post processing methods on generated image whilst preserving the photo-realistic nature of image. Here, we test PALADIN against multiple post processing attacks such as brightness, contrast, saturation, sharpness, horizontal flip, crop and gaussian noise and benchmark against WOUAF. Table 4(a) and 4(b) demonstrates PALADIN robust accuracy and FER respectively against WOUAF on all image post processing employed. PALADIN clearly shows stable accuracy while maintaining low FER. Figure 3 shows accuracy variation versus the strength of the post processing, PALADIN shows stable performance and maintains constant bit accuracy and FER over a range of post processing operations normally used to erase or corrupt fingerprint.
  We also evaluate robustness against JPEG compression, from Table 4(a) and 4(b) we infer that PALADIN outperforms WOUAF when employed with JPEG compression with jpeg quality set to 60. Figure 4 shows accuracy variation versus the jpeg compression strenght, PALADIN shows stable performance and maintains constant bit accuracy and FER over a range of jpeg compression, thus proving if incorporated during training PALADIN generalizes well and outperforms WOUAF.
5 Conclusions
Near-perfect fingerprinting strategies for latent diffusion models, even with 99% attribution accuracies, cannot be deployed due to the unacceptably large absolute number of incorrect predictions involved. We introduce a framework, PALADIN, which preserves generated image quality and also achieves 100% user-attribution accuracy. PALADIN is developed on top of a recent state-of-the-art model, WOUAF. We show how the performance of WOUAF can be further improved by a combination of modified fingerprint decoder architecture and a loss function better equipped to handle the trade-off between generated image quality and the attribution accuracy. Next, we further enhance the model performance by incorporating the ability of bit error recovery, a concept borrowed from the literature of coding theory. Although PALADIN is built atop WOUAF, it is not restricted to this one model. PALADIN provides a plug-and-play layer of cyclic error correction scheme on top of similar models and is capable of making them 100% user-attribution accurate, along with an ability to appropriately flag non-identifiable fingerprints. In this work, we also introduce a new metric FER for measuring fingerprinting accuracy, that is more accurate and informative. To the best of our knowledge, PALADIN is the first ever strategy to achieve 100% user attribution accuracy in text-to-image diffusion models, which makes it possible for real-world deployment.